A point on a rotating turntable 18.5 cm from the center accelerates from rest to a final speed of 0.710 m/s in 1.75 s. At t = 1.21 s, find the magnitude and direction of each of the following. (a) the radial acceleration m/s2 toward the center away from the center in the direction of the rotation in a direction opposite to the rotation (b) the tangential acceleration m/s2 toward the center away from the center in the direction of the rotation in a direction opposite to the rotation (c) the total acceleration of the point m/s2 °
Answer #1

a)

From

V=Vi+at

Tangential acceleration is

=>at=(V-Vi)/t =(0.71-0)/1.75

at = 0.4057 m/s2

Speed at t=1.21 s

V=0+0.4057*1.21 =0.491 m/s

Radial acceleration

ar=V2/r =0.4912/0.185 =1.303 m/s2

b)

Tangential acceleration is

at = 0.4057 m/s2

c)

total acceleration magnitude is

a=sqrt[0.40572+1.3032] =1.3644 m/s2

Direction

o=tan-1(1.303/0.4057) =72.7o (inward from the tangent)